Stock Y has a beta of 1.30 and an expected return of 15.3%. Stock Z has a beta of
0.70 and an expected return of 9.3%. If the risk-free rate if 5.5% and the market risk
premium is 6.8%, are these stocks correctly priced?
Answer #1

In this question we need to compute the reward to risk ratio which is computed as shown below:

Reward to risk ratio of stock Y is computed as follows:

= ( Expected return - risk free rate ) / Beta

= ( 0.153 - 0.055 ) / 1.30

= 0.0754 or 7.54% Approximately

Reward to risk ratio of stock Z is computed as follows:

= ( Expected return - risk free rate ) / Beta

= ( 0.093 - 0.055 ) / 0.7

= 0.0543 or 5.43% Approximately

Reward to risk ratio for the market is

= 0.068 / 1 (Since the beta of the market is 1)

= 0.068 or 6.8%

Since the reward to risk ratio of stock Y is greater than the reward to risk ratio of market, hence the same is undervalued

Since the reward to risk ratio of stock Z is less than the reward to risk ratio of market, hence the same is overvalued

Feel free to ask in case of any query relating to this question